Circular No. 8968

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 Q3 (GARRADD) G. J. Garradd reports his discovery of a comet on CCD images obtained

with the 0.5-m Uppsala Schmidt telescope, the object showing a 10′′ coma (more prominent toward the north); despite poor seeing, the comet looks more diffuse than nearby stars. Following posting on the Minor Planet Center’s ‘NEOCP’, A. C. Gilmore and P. M. Kilmartin write that two stacked 120-s frames taken with the Mt. John 1.0-m reflector on Aug. 28.5 UT show a centrally condensed coma of diameter 6′′ with no tail.

2008 UT  α_{2000}  δ_{2000}  Mag.
Aug. 27.63903  0^{h}29^{m}46.^{s}14  −48^{o}22′19′′.7  19.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-Q52.

T = 2009 July 1.896 TT  ω = 320.^{o}598
Ω = 218.057 }2000.0
q = 2.52970 AU  i = 139.820

V1647 ORIONIS K. Itagaki, Yamagata, Japan, reports (via S. Nakano, Sumoto; and

via H. Yamaoka, Kyushu University) the apparent brightening of V1647 Ori and its associated nebula (cf. IAUCs 8284, 8354, 8396, 8600, 8681, 8694) to mag ∼ 16.7 on an unfiltered CCD image taken on Aug. 26.789 UT with a 0.60-m f/5.7 reflector, noting that neither the variable star nor the associated nebula were visible on his survey frame taken with a 0.21m reflector on Jan. 2.598 (limiting mag 17.5). Itagaki adds that the star appears in the USNO-B1.0 catalogue (epoch 1971.3) at blue mag 19.7 and red mag 19.5, and he provides the following precise position for V1647 Ori from his recent image: α = 5^{h}46^{m}13^{s}.16, δ = −0^{o}06′03′′.6 (equinox 2000.0). Yamaoka notes that a similar quiescent brightness was noted in 2005 Oct. by Acosta-Pulido et al. (2007, A.J. 133, 2020) following its earlier outburst (to mag ∼ 15–16) in 2004 Jan., while an earlier outburst in 1966–1967 was found in the Asiago plate collection by Aspin et al. (2006, A.J. 132, 1298).

COMET 6P/d’ARREST Total-magnitude estimates by J. J. Gonzalez, Leon, Spain: Aug. 5.00

UT, 9.6 (25×100 binoc.); 9.03, 8.9; 23.91, 8.1; 28.00, 8.0 (10×50 binoc.).

2008 August 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green